# What will be the Euclidean dimension of an Ising ferromagnetic cubic shell?

### v1

@@
Abstract: The equilibrium and nonequilibrium properties of an Ising ferromagnetic cubic shell have been extensively studied by Monte Carlo simulation using Metropolis single spin flip algorithm. Although geometrically the Euclidean dimension of the cubical shell is three, interestingly, the Ising ferromagnetic cubic shell undergoes ferromagnetic phase transition at a temperature which is very close to that for two-dimensional Ising ferromagnet. Surprisingly, the Ising ferromagnetic cubic shell shows a strange (neither exponential nor stretched exponential) kind of relaxation behaviour, instead of exponential relaxation as usually observed in the two dimensional Ising ferromagnet. The metastable lifetime of a ferromagnetic Ising cubical shell is studied as a function of the applied magnetic field. Here also, the cubic shell behaves more likely a two-dimensional object as found from statistical analysis and comparison with Becker-Döring prediction of classical nucleation theory.
@@ 1 Introduction:
The equilibrium phase transition of a ferromagnetic sample is well studied[1]. Since, the Ising ferromagnet is a prototype of ferromagnet, the equilibrium ferro-para phase transition is well studied by Monte Carlo simulation. In two dimensions, the model (with nearest neibour interaction) is exactly solvable. In three dimensions, it hase been studied by Monte Carlo simulations. The relaxational behaviours of a ferromagnetic sample is also studied widely. The Ising ferromagnet has been found to show the exponential relaxation both in two and in three dimensions.
@@
The ferromagnetic samples show the metastable behaviour. This metastability is a subject of modern research in the context of preparing the magnetic storage devices. In our everyday life, magnetic thin films are widely used in a rage of devices due to their enormous technological applications. Magnetic data storage [2] are essential and useful devices that play a crucial role in storing information in tiny magnetic grains. The switching of magnetisations by an externally applied magnetic field determines the dynamical behaviours of such storage devices. In fact, the devices for fast recording require rapid switching of magnetisation. For fast recording or rapid accessing the data the magnetic systems are expected to respond quickly to the variations of the external magnetic field. Not only this quick response is sufficient for good storage devices. The magnetic systems should possess adequate stability against any kind of noises (thermal or any non-thermal)[3].
@@
The subject of switching of magnetisation by externally applied weak magnetic field is important to study. Theoretically or historically the problem was investigated by Becker and Döring (in 1935)[4, 5] through their celebrated classical nucleation theory. The reversal time (or nucleation time) has been found phenomenologically as function of temperature and applied magnetic field in general d-dimensions. This picture provides the metastable lifetime and its dependeces on the temperature and applied magnetic field. The rate of nucleation of crystalline solids in a solid-melt system was explored in an important historical study[6].
@@
A considerable amount of theoretical and experimental works has been done over the last ninty years after Becker-Döring theory. The simple and prototypical model of ferromagnet, the Ising ferromagnet, has been chosen to study the reversal phenomena. The relaxation of Ising ferromagnet after a sudden reversal of applied magnetic field is also studied[7]. The dynamics of the domains or droplets of magnetic films with perpendicular anisotropy have been reported[8]. Good agreement between the numerical results and the theoretical predictions of Becker-Döring has been reported[9] through extensive simulational investigations on nucleation in different dimensions using heat-bath dynamics. The prediction of this phenomenological theory was verified by the Monte Carlo simulation where the growth of droplets has been studied as phase ordering kinetics[10]. The rates of growth and decay of the droplets of different sizes have been studied [11] as functions of external field and temperature. The heat-assisted magnetisation reversal in ultrathin films for ultra-high-density information recording has been investigated[12]. The extensivity property of the metastable lifetime ahs also been reported[13]. Very recently, the general S-spin Ising model has been used to study[14] the metastable behaviour and compared the results with Becker-Döring analysis. However, the literature of metastable behaviours in ferromagnetic system is lacking by studies of the role of geometry. Recently, the metastablity can be affected by the geometry of the system. The metastable behaviour of an Ising strips (of equal area) has been studied [15] by Monte Carlo simulation.
@@
What will be the Euclidean dimension of an Ising ferromagnetic cubic shell? Geometrically, it is a three dimensional (Euclidean) object. But will it behave like a three dimensional object in the context of ferromagnetic behaviour ? To address this question, we report, in this paper, the Monte Carlo results of the equilibrium behaviour, relaxational behaviour and the metastable behaviour of an Ising ferromagnet cubic shell. Will it show its tri-dimensionality in the context of ferromagnetic phase transition ? Will it show its tri-dimensionality in the context of magnetic relaxation behaviour? Will it show its tri-dimensionality in the context of growth of droplets in the reversal of magnetisation? All these questions are addressed (through extensive Monte Carlo investigation) in this article. The manuscript is organized as follows: the model of the cubical shell of Ising ferromagnet alongwith the simulational scheme are mentioned in the next section, the numerical results are provided in section-3, the paper ends with summary in section-4.
@@ 2 Model and Simulation:
The Monte Carlo imulations have been executed on a cubic shell of size $L$, having $N=L^{3}-(L-2)^{3}=6L^{2}-12L+8$ number of lattice sites with open boundary conditions in all three directions. As a simple demonstration, a schematic diagram of a cubic shell of $L=4$ is shown in figure 1. The blue coloured balls denote the spins on the suface of the cubic shell lattice, while red coloured balls indicate the spins on the edge or boundary of the lattice. Inside the cube, there is no such magnetic spins. At any specified temperature ($T$), the probability of spin-flip at any site $i$ (randomly chosen) is determined by Metropolis algorithm[16].
@@ 3.1 Equilibrium behaviour:
In this study, we have investigated the temperature dependence of magnetisation and susceptibility in the Ising ferromagnetic system modelled on a cubic shell lattice of $L=100$. The thermodynamic quantities ($M~{}\rm and~{}\chi$) are recorded up to 75000 MCSS time scale, out of which the initial 65000 MCSS time-steps are discarded to achieve equilibrium. These statistical quantities ($M~{}\rm and~{}\chi$) are determined by taking the average over the last $10^{4}$ MCSS. Furthermore, we have calculated the sample average of $M~{}\rm and~{}\chi$ over 250 random samples. We have cooled down the system with small step size ($\delta T$) of the temperature. The step size is chosen $\Delta T=0.02$ near the transition temperature. However, larger step size ($\Delta T=0.1$) has been chosen for the temperature much away from the critical temperature. To determine the transition point, the temperature dependence of both magnetisation $M$ and susceptibility $\chi$ are studied and graphically depicted in figure 2(a) and 2(b). The average magnetisation ($M$) continuously grows from zero to nonzero, as the system cools down and the peak of susceptibility ($\chi$) at a particular temperature $T=2.28$. This indicates the continuous ferro-para phase transition of an Ising ferromagnetic cubic shell. This transition point is defined as the pseudo-critical temperature of the Ising ferromagnetic cubic shell, i.e., $T_{c}^{p}=2.28$ (in unit of $J/k_{B}$) for the finite-sized cubic shell of $L=100$. At thermodynamic limit ($L\rightarrow\infty$) the value critical temperature $T_{c}$ for Ising cubic shell should be slightly different from $T_{c}^{p}$. However, it is worth mentioning that, the value of $T_{c}^{p}$ for the cubic shell is remarkably close to the Onsager value of $T_{c}=2.269...$[17] in unit of $J/k_{B}$ for the two-dimensional Ising system. In the respect of ferromagnetic phase transition, the Ising ferromagnetic cubic shell behaves like a two-dimensional object. This result is not at all surprising. As the Ising ferromagnetic cubic shell is an object where each spin (except the spins at eight corners) has four nearest neighbours, it is not surprising to find the behaviour of a two dimensional object in the context of equilibrium ferro-para phase transition.
@@
We have also observed the phase transition in the solid cubic Ising ferromagnetic system of lattice size $L=50$ with open boundary condition, analysing the variation of average magnetisation and susceptibility with temperature. The pseudo-critical temperature for solid cubic lattice is marked to be $T_{c}^{p}=4.46$ (figure 2 (c) and (d)), which is close to the critical point of a three dimensional system $T_{c}=4.511$ [18]. The comparative study of the transition point between a cubic shell and a Solid cube is demonstrated in figure 2. So, we have come to the conclusion that the ferro-para phase transition in Ising ferromagnetic cubic shell matches that for a two-dimensional system, even though geometrically both cubic shell or hollow cube are three-dimensional object. Should we expect to have the two dimensional behaviour of an Ising ferromagnetic cubic shell in the context of relaxational behaviour? To get the answer to this question, we have studied the relaxation behaviour of an Ising ferromagnetic cubic shell and compared the results with that for a solid cube.
@@
Since we are eager to examine transient behaviours like magnetic relaxation, and metastable dynamics of Ising ferromagnetic cubic shell along with a comparative study between cubic shell and solid cube, the equivalent thermal condition should be maintained in both cases. For that reason, we have chosen the temperature of the respective systems as with same multiplicative factor of respective critical temperature $T_{c}$. Precisely, if the temperature for the Ising ferromagnetic cubic shell is $T=fT_{c}^{Shell}$, the same thermal condition will be provided for solid cube at temperature $T=fT_{c}^{solid}$. In the next subsection, we will discuss the magnetic relaxation behaviour of both Ising ferromagnetic cubic shell and solid cube.
@@ 3.2 Relaxation behaviour:
The ferromagnetic model system exhibits well studied relaxation behaviours. If the system is kept initially at any cofiguration of all spins up ($S_{i}=+1$) for any finite temperature $T$. The system will relax to eventually reaches its equilibrium state for that temperature. Here, we have studied such kind of relaxation behaviour of an Ising ferromagnetic cubic shell. Consider an Ising ferromagnetic cubic shell (of size $L=50$) with Ising ferromagnetic spins, initially $S_{i}=+1\forall i$. Certainly, this is not the equilibrium state of the system in its respective paramagnetic phase. We are now trying to have its equilibrium state for a temperature slightly above the corresponding critical temperature. Now let the system relax in the absence of an external magnetic field ($h_{ext}=0$). The magnetisation of the system $m(t)$ starts to decay with time and eventually vanishes. This relaxation process is not instantaneous; it takes a certain time duration and finally, the system achieves the stable equilibrium state of $m=0$.
@@
We have investigated the magnetic relaxation in the Ising ferromagnetic cubic shell for three different temperatures (fixed at $T=1.05T_{c}^{p},~{}1.08T_{c}^{p}$ and $1.10T_{c}^{p}$), as depicted in figure 3 (a) and (b). The instantaneous magnetisation $m(t)$ is determined by taking average over 10000 random samples to have the smooth variation with time. We have studied the instantaneous magnetisation ( sample averaged) as a function of time ($t$) and plotted in both linear and semi-logarithmic scale for cubical shell as well as a solid cubic lattice of size $L=50$ with open boundary condition. Figure 3 (a) and (b) illustrate the time evolution of magnetisation for cubic shell lattice in linear and semi-log scale respectively, while figure 3 (c) and (d) represent those for the solid cube at three different temperatures. The theoretical [19] study as well as Monte Carlo simulation predicted[15] the exponential nature of magnetic relaxation both in two and three-dimensional Ising ferromagnetic systems near critical point. The variation of magnetisation with time in semi-log scale (Fig-3(d)) gives a straight line for solid cubic lattice, revealing the exponential relaxation. Whereas our results for cubic shell exhibit a non-exponential relation in the semi-log plot of $m(t)~{}vs.~{}t$(Fig-3(b)). Here is the surprise !! We have seen (in the previous subsection) that the equilibrium ferro-para phase transition of an Ising ferromagnetic cubic shell compels us to believe that it behaves like a two dimensional object. But the two dimensional Ising ferromagnet shows the exponential behaviours as contradicted by our present result of non-exponential relaxation of an Ising ferromagnetic cubic shell.
@@
So, what kind of relaxation behaviour is expected from such a cubic shell ? Is it stretched exponential ? Is it power law type ? we have examined the variation of $\log(\log(m(t))$ with time to ascertain whether the cubic shell undergoes stretched exponential relaxation or not. Fig-4 shows the plot of $\log(\log(100\times m(t))$ as a function of time ($t$). The results indicate that magnetic relaxation of the Ising ferromagnetic cubical shell does not even show the stretched exponential behaviour. The logarithmic value of magnetisation $\log(m(t))$ is also studied as a function of $\log(t)$ and shown in Fig-5. This graph also fails to indicate any power-law type of relaxation behaviour. May we conclude from these results of relaxation that the Ising ferromagnetic cubic shell even does not behave like a two dimensional ojbect as predicted from its equilibrium ferro-para phase transition?
@@
The strange behaviours of the cubic shell prompted us to study further its metastable behaviour and compare that with the prediction of classical nucleation theory. The classical nucleation theory (Becker-Döring theory) systematically and critically provides the role of Euclidean dimensions of the ferromagnetic system through its dependence reflected in the magnetisation reversal time. The next subsection is devoted to the Monte Carlo study of metastable behaviours of an Ising ferromagnetic cubic shell and that of a solid cubic Ising ferromagnet.

### v2

@@
Abstract: The equilibrium and nonequilibrium properties of an Ising ferromagnetic cubic shell have been extensively studied by Monte Carlo simulation using Metropolis single spin flip algorithm. Although, geometrically the Euclidean dimension of the cubical shell is three, interestingly, the Ising ferromagnetic cubic shell undergoes ferromagnetic phase transition at a temperature which is very close to that for two-dimensional Ising ferromagnet. Surprisingly, the Ising ferromagnetic cubic shell shows a strange (neither exponential nor stretched exponential) kind of relaxation behaviour, instead of exponential relaxation as usually observed in the two dimensional Ising ferromagnet. The metastable lifetime of a ferromagnetic Ising cubical shell is studied as a function of the applied magnetic field. Here also, the cubic shell behaves more likely a two-dimensional object as found from statistical analysis and comparison with Becker-Döring prediction of classical nucleation theory.
@@ 1 Introduction:
The equilibrium phase transition of a ferromagnetic sample is well studied[1]. Since, the Ising ferromagnet is a prototype of ferromagnet, the equilibrium ferro-para phase transition is well studied by Monte Carlo simulation. In two-dimensions, the model (with nearest neibour interaction) is exactly solvable. In three-dimensions, it hase been studied by Monte Carlo simulations. The relaxational behaviours of a ferromagnetic sample is also studied widely. The Ising ferromagnet has been found to show the exponential relaxation both in two and in three dimensions.
@@
The ferromagnetic samples show the metastable behaviour. This metastability is a subject of modern research in the context of preparing the magnetic storage devices. In our everyday life, magnetic thin films are widely used in a range of devices due to their enormous technological applications. Magnetic data storages [2] are essential and useful devices that play a crucial role in storing information in tiny magnetic grains. The switching of magnetisations by an externally applied magnetic field determines the dynamical behaviours of such storage devices. In fact, the devices for fast recording require rapid switching of magnetisation. For fast recording or rapid accessing the data the magnetic systems are expected to respond quickly to the variations of the external magnetic field. Not only this quick response is sufficient for good storage devices. The magnetic systems should possess adequate stability against any kind of noise (thermal or any non-thermal)[3].
@@
The subject of switching of magnetisation by externally applied weak magnetic field is important to study. Theoretically or historically the problem was investigated by Becker and Döring (in 1935)[4, 5] through their celebrated classical nucleation theory. The reversal time (or nucleation time) has been found phenomenologically as a function of temperature and applied magnetic field in general d-dimensions. This picture provides the metastable lifetime and its dependences on the temperature and applied magnetic field. The rate of nucleation of crystalline solids in a solid-melt system was explored in an important historical study[6].
@@
A considerable amount of theoretical and experimental works has been done over the last ninety years after Becker-Döring theory. The simple and prototypical model of ferromagnet, the Ising ferromagnet, has been chosen to study the reversal phenomena. The relaxation of Ising ferromagnet after a sudden reversal of applied magnetic field is also studied[7]. The dynamics of the domains or droplets of magnetic films with perpendicular anisotropy have been reported[8]. Good agreement between the numerical results and the theoretical predictions of Becker-Döring has been reported[9] through extensive simulational investigations on nucleation in different dimensions using heat-bath dynamics. The prediction of this phenomenological theory was verified by the Monte Carlo simulation where the growth of droplets has been studied as phase ordering kinetics[10]. The rates of growth and decay of the droplets of different sizes have been studied [11] as functions of external field and temperature. The heat-assisted magnetisation reversal in ultrathin films for ultra-high-density information recording has been investigated[12]. The extensivity property of the metastable lifetime has also been reported[13]. Very recently, the general S-spin Ising model has been used to study[14] the metastable behaviour and compared the results with Becker-Döring analysis. However, the literature of metastable behaviours in ferromagnetic system is lacking by studies of the role of geometry. Recently, the metastablity can be affected by the geometry of the system. The relaxation behaviour of an Ising strips (of equal area) has been studied [15] by Monte Carlo simulation.
@@
What will be the Euclidean dimension of an Ising ferromagnetic cubic shell? Geometrically, it is a three dimensional (Euclidean) object. But will it behave like a three dimensional object in the context of ferromagnetic behaviour ? To address this question, we report, in this paper, the Monte Carlo results of the equilibrium behaviour, relaxational behaviour and the metastable behaviour of an Ising ferromagnetic cubic shell. Will it show its tri-dimensionality in the context of ferromagnetic phase transition ? Will it exhibit its tri-dimensionality in the context of magnetic relaxation behaviour? Will it reveal its tri-dimensionality in the context of growth of droplets in the reversal of magnetisation? All these questions are addressed (through extensive Monte Carlo investigation) in this article. The manuscript is organized as follows: the model of the cubical shell of Ising ferromagnet along with the simulational scheme are mentioned in the next section, the numerical results are provided in section-3, the paper ends with summary in section-4.
@@ 2 Model and Simulation:
The Monte Carlo simulations have been executed on a cubic shell of size $L$, having $N_{s}=L^{3}-(L-2)^{3}=6L^{2}-12L+8$ number of lattice sites with open boundary conditions in all three directions. The reason, for the implementation of open boundary condition here, is as follows: the spin of each surface will interact with the neighbouring spins of that surface only except from the spins at the edges. The spins at the edges will interact with neighbouring spins of two adjuscent surfaces. The periodic boundary condition, if implemented, would provide additional interaction for the spins located on any surface of the cubical shell.
As a simple demonstration, a schematic diagram of a cubic shell of $L=4$ is shown in figure 1. The blue coloured balls denote the spins on the suface of the cubic shell lattice, while red coloured balls indicate the spins on the edge or boundary of the lattice. Inside the cube, there is no such magnetic spins. At any specified temperature ($T$), the probability of spin-flip at any site $i$ (randomly chosen) is determined by Metropolis algorithm[16].
@@ 3.1 Equilibrium behaviour:
In this study, we have investigated the temperature dependence of magnetisation and susceptibility in the Ising ferromagnetic system modelled on a cubic shell of size $L=100$. The thermodynamic quantities ($M~{}\rm and~{}\chi$) are recorded up to 75000 MCSS time steps, out of which the initial 65000 MCSS time-steps are discarded to achieve equilibrium. These statistical quantities ($M~{}\rm and~{}\chi$) are determined by taking the average over the last $10^{4}$ MCSS. Furthermore, we have calculated the sample average of $M~{}\rm and~{}\chi$ over 250 random samples. We have cooled down the system with small step size ($\delta T$) of the temperature. The step size is chosen $\Delta T=0.02$ near the transition temperature. However, larger step size ($\Delta T=0.1$) has been chosen for the temperature much away from the critical temperature. To determine the transition point, the temperature dependences of both magnetisation $M$ and susceptibility $\chi$ are studied and graphically depicted in figure 2(a) and 2(b). The average magnetisation ($M$) has been found to grow continuously from zero to nonzero, as the system cools down and the susceptibility ($\chi$) has been found to get peaked at a particular temperature $T=2.28$. This indicates the continuous ferro-para phase transition of an Ising ferromagnetic cubic shell. This transition point is defined as the pseudo-critical temperature of the Ising ferromagnetic cubic shell, i.e., $T_{c}^{p}=2.28$ (in unit of $J/k_{B}$) for the finite-sized ($L$ finite) cubic shell of $L=100$. In the thermodynamic limit ($L\rightarrow\infty$) the value of the critical temperature $T_{c}$ for Ising cubic shell should be slightly different from $T_{c}^{p}$. However, it is worth mentioning that, the value of $T_{c}^{p}$ for the cubic shell is remarkably close to the Onsager value of $T_{c}=2.269...$[17] (in unit of $J/k_{B}$) for the two-dimensional Ising system. In the context of ferromagnetic phase transition, the Ising ferromagnetic cubic shell behaves here like a two-dimensional object. This result is not at all surprising. Since the Ising ferromagnetic cubic shell is a structure where each spin (except the spins at eight corners) has four nearest neighbours, its behaviour regarding the equilibrium ferro-para phase transition resembles that of a two-dimensional system.
@@
We have also observed the phase transition in the solid cubic Ising ferromagnetic system of lattice size $L=50$ with open boundary condition, analysing the variation of average magnetisation and susceptibility with temperature. The pseudo-critical temperature for solid cubic lattice is marked to be $T_{c}^{p}=4.46$ (figure 2 (c) and (d)), which is close to the Monte Carlo estimate of the critical temperature of a three dimensional system $T_{c}=4.511$ [18]. The comparative study of the transition point between a cubic shell and a solid cube is demonstrated in figure 2. So, we have come to the conclusion that the ferro-para phase transition in Ising ferromagnetic cubic shell matches with that for a flat two-dimensional system, even though geometrically both cubic shell and solid cube are three-dimensional objects. Should we expect to have the two dimensional behaviour of an Ising ferromagnetic cubic shell in the context of relaxational behaviour? To find the answer to this question, we have studied the relaxation behaviour of an Ising ferromagnetic cubic shell and compared the results with that for a solid cube.
@@
Since we are eager to examine transient behaviours like magnetic relaxation, and metastable dynamics of Ising ferromagnetic cubic shell along with a comparative study between cubic shell and solid cube, the equivalent thermal condition should be maintained in both cases. For that reason, we have chosen the temperature of the respective systems with same multiplicative factor of respective critical temperatures ($T_{c}$). Precisely, if the temperature for the Ising ferromagnetic cubic shell is $T=f\times T_{c}^{Shell}$, the same thermal condition will be provided for solid cube at temperature $T=f\times T_{c}^{solid}$. In the next subsection, we will discuss the magnetic relaxation behaviour of both Ising ferromagnetic cubic shell and solid cube.
@@ 3.2 Relaxation behaviour:
The ferromagnetic model system exhibits well studied relaxation behaviours. If the system is initially configured with all spins up ($S_{i}=+1~{}\forall~{}i$) at any finite temperature T, it will eventually relax and reach its equilibrium state corresponding to that temperature ($T$). Here, we have studied such kind of relaxation behaviour of an Ising ferromagnetic cubic shell. Consider an Ising ferromagnetic cubic shell (of size $L=50$), initially $S_{i}=+1~{}\forall~{}i$. Certainly, this is not the equilibrium state of the system at any finite nonzero temperature $T$. We are trying to achieve its equilibrium state for a temperature slightly above the corresponding critical temperature. Now let the system relax in the absence of an external magnetic field ($h_{ext}=0$). The magnetisation of the system $m(t)$ starts to decay with time and eventually vanishes. This relaxation process is not instantaneous; it takes a certain time duration and finally, the system achieves the stable equilibrium state of $m=0$.
@@
We have investigated the magnetic relaxation in the Ising ferromagnetic cubic shell for three different temperatures (fixed at $T=1.05T_{c}^{p},~{}1.08T_{c}^{p}$ and $1.10T_{c}^{p}$), as depicted in figure 3 (a) and (b). The instantaneous magnetisation $m(t)$ is determined by taking average over 10000 random samples to ensure the smooth variation with time. We have studied the instantaneous magnetisation ( sample averaged) as a function of time ($t$) and plotted in both linear and semi-logarithmic scale for cubical shell as well as a solid cubic lattice of size $L=50$ with open boundary condition. Figure 3 (a) and (b) illustrate the time evolution of magnetisation for cubic shell in linear and semi-log scale respectively, while figure 3 (c) and (d) represent those for the solid cube at three different temperatures. The theoretical [19] study as well as Monte Carlo simulation predicted[15] the exponential nature of magnetic relaxation both in two and three-dimensional Ising ferromagnetic systems near critical point. The variation of magnetisation with time in semi-log scale (Fig-3(d)) gives a straight line for solid cubic lattice, revealing the exponential relaxation. Whereas our results for cubic shell exhibit a non-exponential relaxation in the semi-log plot (not linear) of $m(t)$  versus $t$ (Fig-3(b)). Here is the surprise !! We have seen (in the previous subsection) that the equilibrium ferro-para phase transition of an Ising ferromagnetic cubic shell compels us to believe that it behaves magnetically like a two-dimensional system. But the two dimensional Ising ferromagnet shows the exponential behaviours as contradicted by our present result of non-exponential relaxation of an Ising ferromagnetic cubic shell. Usualy, exponential relaxational is observed in two-dimensional Ising ferromagnet and that is theoretically understood in the linearized mean field equation of Glauber kinetic Ising ferromagnet[19]. Those results are obtained for the compact solid systems. However, the studies on relaxational behaviours for the hollow systems (or fractal like systems) are missing in the literature. The systems (not simply connected) with large cavity may show different relaxtion behaviours. It is not yet understood the basic reason of having non-exponential relaxation in the case of cubic shell of Ising ferromagnet. The results of our present study may trigger further research to investigate the basic reasons of magnetic relaxation related to the geometry (not simply connected) of the system. It may be worth mentioning here that the non-exponential magnetic relaxation has recently been observed experimentally in magnetic nanoparticles [20]. The single molecular magnet of manganese also showed non-exponential magnetic relaxation[21].
@@
So, what kind of relaxation behaviour is expected from such a cubic shell ? Is it stretched exponential ? Is it power law type ? We have examined the variation of $\log(\log(m(t))$ with time to ascertain whether the cubic shell undergoes stretched exponential relaxation or not. Fig-4 shows the plot of $\log(\log(100\times m(t))$ as a function of time ($t$). The results indicate that magnetic relaxation of the Ising ferromagnetic cubical shell does not even show the stretched exponential behaviour. The logarithmic value of magnetisation $\log(m(t))$ is also studied as a function of $\log(t)$ and shown in Fig-5. This graph also fails to indicate any power-law type of relaxation behaviour. May we conclude from the obtained results of the non-exponential relaxation behaviour that the Ising ferromagnetic cubic shell does not even behave like a geometrically two-dimensional object, as predicted from its equilibrium ferro-para phase transition (section-3.1)? So, what will be the Euclidean dimension of an Ising ferromagnetic cubic shell?
@@
The strange relaxation behaviours of the cubic shell prompted us to proceed further to investigate its metastable behaviour and compare that with the prediction of classical nucleation theory. The classical nucleation theory (Becker-Döring theory) systematically and critically provides the role of Euclidean dimensions of the ferromagnetic system through its (dimensional) dependence reflected in the magnetisation reversal time. The next subsection is devoted to the Monte Carlo study of metastable behaviours of an Ising ferromagnetic cubic shell and that of a solid cubic Ising ferromagnet.